### v1

# Fractional binomial regression model for count data with excess zeros

Abstract.
This paper proposes a new generalized linear model with fractional binomial distribution. Zero-inflated Poisson/negative binomial distributions are used for count data that has many zeros. To analyze the association of such a count variable with covariates, zero-inflated Poisson/negative binomial regression models are widely used. In this work, we develop a regression model with the fractional binomial distribution that can serve as an additional tool for modeling the count response variable with covariates. Data analysis results show that on some occasions, our model outperforms the existing zero-inflated regression models.
Email address: C. Breece: ceb1386@uncw.edu, J. Lee (corresponding author): leejb@uncw.edu.
Keywords: Zero-inflated regression models, count data, fractional binomial distribution.
1. Introduction
Count data with excess zeros are often found in many areas, including public health, epidemiology, ecology, finance, and quality control, to name a few. In the past, several zero-adjusted discrete models have been developed for such a data set [3]. When there is additional information from covariates and one wishes to analyze the association of covariates with a count response variable with many zeros, then the zero-inflated Poisson/negative binomial regression model can be used [2, 6, 7, 9]. In [4], the zero-inflated binomial regression model was developed with random effects that can account for correlated responses and multiple sources of variance. These zero-inflated models are mixed models in which structural zeros are modeled separately and added to a regular count model that generates sampling zeros and other count values.
In this work, we propose a new model that can serve as a regression model for count data with excess zeros. This model, which we will call a fractional binomial regression model, is developed from the fractional binomial distribution first introduced in [8]. It does not treat sampling zeros and structural zeros separately, yet it is flexible as much as or more than the conventional zero-inflated models so that it can fit count data of various shapes.
The fractional binomial distribution is defined by the sum of dependent Bernoulli variables, more specifically, the sum in a stationary binary sequence whose covariance function decays by power law. The mean of the fractional binomial distribution is the same as that of the regular binomial distribution since it is defined from a stationary sequence, but the variance of the fractional distribution is larger than that of the regular binomial distribution due to the dependence among the Bernoulli variables. Especially, under certain conditions, its variance is proportional to the length of the binary sequence to a fractional power.
The shape of fractional binomial distribution varies with parameters that affect the center, spread, skewness, and zero-inflation of the distribution. Using the logit link with the parameters, one can incorporate covariates into the distribution to associate the covariates with the shape of the distribution of the response variable.
In Section 2, we briefly review the zero-inflated Poisson/negative binomial regression models. In Section 3, we introduce the fractional binomial regression model and investigate the consistency of the ML estimators, followed by simulations to examine the performance of the estimators in Section 4. In Section 5, we use two datasets to fit the zero-inflated regression models and fractional binomial regression model, compare their fit, and test to see whether the fractional binomial regression model is significantly closer to the actual generating process. The conclusion and discussion are followed in Section 6.
2. Zero-inflated regression models
We first review widely used zero-inflated regression models: zero-inflated Poisson and negative regression models. Let $y_{i}\in\mathbb{N}\cup\{0\}$ be a response variable and $(x_{1i},x_{2i},\cdots,x_{ki})\in\mathbb{R}^{k}$ be covariates for $i=1,2,\cdots,n$. Let ${\bf x}_{i}=(1,x_{1i},x_{2i},\cdots,x_{ki}),$ ${\boldsymbol{\gamma}}=(\gamma_{0},\gamma_{1},\gamma_{2},\cdots,\gamma_{k}),$ ${\boldsymbol{\beta}}=(\beta_{0},\beta_{1},\beta_{2},\cdots,\beta_{k}).$ The probability distributions of the Zero-inflated Poisson (ZIP) and zero-inflated negative binomial (ZINB) regression models are expressed as
$P(y_{i}|{\bf x}_{i},\Theta)=\begin{dcases}\pi_{i}+(1-\pi_{i})g(0|{\bf x}_{i})&\text{if }y_{i}=0,\\ (1-\pi_{i})g(y_{i}|{\bf x}_{i})&\text{if }y_{i}\in\mathbb{N},\end{dcases}$
where
$\pi_{i}=\frac{1}{1+\exp(-{\bf x}_{i}^{\prime}{\boldsymbol{\gamma}})},$
$g(y_{i}|{\bf x}_{i})=\begin{dcases}\frac{e^{-\mu_{i}}\mu_{i}^{y_{i}}}{y_{i}!}&\text{for ZIP regression model,}\\ \frac{\theta^{\theta}\mu_{i}^{y_{i}}\Gamma(\theta+y_{i})}{\Gamma(\theta)\Gamma(y_{i}+1)(\theta+\mu_{i})^{\theta+y_{i}}}&\text{for ZINB regression model,}\end{dcases}$
and
$\mu_{i}=\exp({\bf x}_{i}^{\prime}{\boldsymbol{\beta}}).$
Let $\Theta$ denote the set of all parameters in the corresponding model, i.e., in ZIP, $\Theta=({\boldsymbol{\beta}},{\boldsymbol{\gamma}}),$ and in ZINB-2, $\Theta=({\boldsymbol{\beta}},{\boldsymbol{\gamma}},\theta).$ Note that $g$ is the probability distribution of the Poisson/negative binomial regression model where $\mu_{i}$ is the conditional mean given the covariates, and $\pi_{i}$ gives the additional probability at zero.
In the Poisson regression model, $y_{i}|{\bf x}_{i}\sim$ Poisson($\mu_{i}$), given the values of covariates, the conditional mean and the conditional variance are the same as $\mu_{i}.$ In the negative binomial regression model, $y_{i}|{\bf x}_{i}\sim$NB($\mu_{i}$), the conditional variance is $\mu_{i}+\mu_{i}^{2}/\theta,$ therefore, the negative binomial distribution can serve for overdispersed count data.
We can also generalize the dispersion parameter $\theta$ in ZINB and replace $\theta$ by
$\theta_{i}=\exp({\bf x}_{i}^{\prime}{\boldsymbol{\alpha}}),$
where $\boldsymbol{\alpha}=(\alpha_{0},\alpha_{1},\cdots,\alpha_{k})$, and we call it ZINB-2.
The parameters of these models are estimated by the maximum likelihood estimation method,
$\hat{\Theta}=\operatorname*{arg\,max}_{\Theta\in\Omega}L(\Theta|(y_{i},{\bf x}_{i}),i=1,2,\cdots,n)=\operatorname*{arg\,max}_{\Theta\in\Omega}\prod_{i=1}^{n}P(y_{i}|{\bf x}_{i},\Theta)$
where $\Omega=\mathbb{R}^{dim(\Theta)}.$ The asymptotic properties of the ML estimators of ZIP was studied in [1], and the ML and the quasi-likelihood estimation of zero-inflated models are discussed in [12].
3. Fractional binomial regression model
Fractional binomial distribution was introduced in [8] and it is defined by the cumulative sum in a sequence of dependent Bernoulli trials, called a generalized Bernoulli process (GBP). The GBP, $\{\xi_{i},i\in\mathbb{N}\}$, was defined as follows.
$P(\xi_{i}=1)=p,P(\xi_{i}=0)=1-p,$
$P(\xi_{i_{0}}=1,\xi_{i_{1}}=1,\cdots,\xi_{i_{n}}=1)=p\prod_{j=1}^{n}(p+c|i_{j}-i_{j-1}|^{2H-2}),$ (3.1)
and for any disjoint sets $A,B\subset\mathbb{N},$
$P(\cap_{i^{\prime}\in B}\{\xi_{i^{\prime}}=0\}\cap_{i\in A}\{\xi_{i}=1\})=\sum_{k=0}^{|B|}\sum_{\begin{subarray}{c}B^{\prime}\subset B\\ |B^{\prime}|=k\end{subarray}}(-1)^{k}P(\cap_{i\in B^{\prime}\cup A}\{\xi_{i}=1\}),$ (3.2)
and
$P(\cap_{i^{\prime}\in B}\{\xi_{i^{\prime}}=0\})$ $=1+\sum_{k=1}^{|B|}\sum_{\begin{subarray}{c}B^{\prime}\subset B\\ |B^{\prime}|=k\end{subarray}}(-1)^{k}P(\cap_{i\in B^{\prime}}\{\xi_{i}=1\}).$ (3.3)
with parameters, $p,H,c$ that satisfy the following assumption.
Assumption 3.1.
$p,H\in(0,1),$ and
$0\leq c<\min\{1-p,\frac{1}{2}(-2p+2^{2H-2}+\sqrt{4p-p2^{2H}+2^{4H-4}})\}.$
The GBP is a stationary binary sequence with a covariance function
${Cov}(\xi_{i},\xi_{j})=pc|i-j|^{2H-2},i\neq j.$
When $H\in(0.5,1),$ the GBP possesses long-range dependence, since $\sum_{i=1}^{\infty}{Cov}(\xi_{1},\xi_{i})=\infty.$ Fractional binomial random variable, denoted by $B_{N}(p,H,c)$, is defined as the sum of the first $N$ variables in the GBP, the mean $E(B_{N}(p,H,c))$ is $Np$, and the variance is
$E((B_{n}-np)^{2})$$\sim\begin{dcases}b_{1}N&\text{ if $H\in(0,.5),$}\\ b_{2}N\ln{N}&\text{ if $H=.5,$}\\ b_{3}N^{2H}&\text{ if $H\in(.5,1),$}\end{dcases}$
where $b_{1}=p(1-p)+2pc/(1-2H),b_{2}=2pc,$ and $b_{3}={pc}/{(H(2H-1))}.$
Note that if $H\in(0.5,1),$ its variance is asymptotically proportional to $N$ to a fractional power. When $c=0,$ $B_{N}(p,H,0)$ becomes the regular binomial random variable whose parameters are $N,p.$
The pmf of the fractional binomial distribution (FB) does not have a closed-form expression, however, it can be computed iteratively through equations (3.1-3.3).
Figure 1 shows the pmfs with various sets of parameters. R package “frbinom” was used to generate the pmfs with $N=30$.
Figure 1. Probability mass function of fractional binomial distribution with changing parameter of p(top), H(middle), c(bottom)
[…]
By connecting the parameters with covariates through a logit link, we can define the fractional binomial regression model as follows. Let ${\boldsymbol{\psi}}=(\psi_{0},\psi_{1},\cdots,\psi_{k}),$ ${\boldsymbol{\eta}}=(\eta_{0},\eta_{1},\cdots,\eta_{k}),$ and ${\boldsymbol{u}}=(u_{0},u_{1},\cdots,u_{k}).$ Given covariates ${\bf x}_{i}\in\mathbb{R}^{k}$, the response variable follows the fractional binomial distribution,
$y_{i}|{\bf x}_{i}\sim B_{N}(p_{i},H_{i},c_{i})$
where the covariates are linked through the logit function of the parameters,
$p_{i}=$ $\frac{1}{1+\exp(-{\bf x}_{i}^{\prime}{\boldsymbol{\psi}})},$ (3.4)
$H_{i}=$ $\frac{1}{1+\exp(-{\bf x}_{i}^{\prime}{\boldsymbol{\eta}})},$ (3.5)
$c_{i}=$ $\frac{1}{1+\exp(-{\bf x}_{i}^{\prime}{\boldsymbol{u}})},$ (3.6)
and $\Theta=({\boldsymbol{\psi}},{\boldsymbol{\eta}},{\boldsymbol{u}})\in\Omega_{\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}\}}$ where $\Omega_{\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}\}}$ is a set of all $\Theta\in\mathbb{R}^{3(k+1)}$ such that $(p_{i},H_{i},c_{i})$ in (3.4-3.6) satisfies Assumption 3.1 for all $i=1,2,\cdots,n.$ One can set $N$ as the maximum value that the response variable can take if it has an upper bound. But in practice, if the sample size is large enough, one can consider the maximum value of the response in data as an approximate upper bound, and let $N=\max\{y_{1},y_{2},\cdots,y_{n}\}.$
The parameters are estimated by the maximum likelihood estimation method,
$\hat{\Theta}=\operatorname*{arg\,max}_{\Theta\in\Omega_{\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}\}}}L(\Theta|(y_{i},{\bf x}_{i}),i=1,2,\cdots,n)=\operatorname*{arg\,max}_{\Theta\in\Omega_{\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}\}}}\prod_{i=1}^{n}P(y_{i}|{\bf x}_{i},\Theta)$
where $P(y_{i}|{\bf x}_{i},\Theta)=P(B_{N}(p_{i},H_{i},c_{i})=y_{i}).$ As the pmf of the fractional binomial distribution does not have an explicit expression, MLE is obtained numerically.
We can reparameterize $(p,H,c)$ in fractional binomial distribution by $(p,H,c^{\circ}),$ where
$c=c^{\circ}\times\min\{1-p,\frac{1}{2}(-2p+2^{2H-2}+\sqrt{4p-p2^{2H}+2^{4H-4}})\},$
so that Assumption 3.1 is satisfied with $(p_{i},H_{i},c_{i}^{\circ})\in(0,1)^{3}.$ As a result, the estimation procedure becomes unconstrained optimization with the parameter space $\Omega=\mathbb{R}^{3(k+1)},$ and
$\hat{\Theta}=\operatorname*{arg\,max}_{\Theta\in\Omega}L(\Theta|(y_{i},{\bf x}_{i}),i=1,2,\cdots,n)=\operatorname*{arg\,max}_{\Theta\in\Omega}\prod_{i=1}^{n}P(y_{i}|{\bf x}_{i},\Theta)$
where $\Theta=({\boldsymbol{\psi}},{\boldsymbol{\eta}},{\boldsymbol{\nu}}),P(y_{i}|{\bf x}_{i},\Theta)=P(B_{N}(p_{i},H_{i},c_{i}^{\circ})=y_{i}),$ and $p_{i},H_{i},c_{i}^{\circ}$ are (3.4), (3.5), and
$c_{i}^{\circ}=\frac{1}{1+\exp(-x_{i}^{\prime}\nu)},$
respectively. As the reparameterization simplifies the parameter space, we will use this version from now on.
For the ML estimator to be consistent in a fractional binomial regression model, we need the following assumptions on the covariates and parameters.
Assumption 3.2.
i) Covariates are i.i.d. random vectors and bounded, i.e., ${\bf X}_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}g$ and the support set of the distribution $g$ is bounded, ${\bf X}_{i}=(x_{i1},x_{i2},\cdots,x_{ik})\in\mathcal{X}$ for all $i=1,2,\cdots$, where $\mathcal{X}$ is a compact set in $\mathbb{R}^{k}$.
ii) Covariates are linearly independent and not constant, i.e., there is no set of constants $(a_{1},a_{2},\cdots,a_{k})\in\mathbb{R}^{k+1}/\{{\bf 0}\}$ such that $var(a_{1}x_{i1}+a_{2}x_{i2}+\cdots+a_{k}x_{ik})=0.$
iii) The true parameter value $\Theta_{0}$ lies in a known compact set, i.e., $\Theta_{0}\in\mathcal{C}$ where $\mathcal{C}$ is a compact set in $\mathbb{R}^{3(k+1)}.$
Here we assume $N$ is known, but if $N$ is not unknown but lies in a known compact set, then one can extend the parameter as $\Theta=({\boldsymbol{\psi}},{\boldsymbol{\eta}},{\boldsymbol{\nu}},N)$ and one can derive the consistency of ML estimator similarly as in the next theorem.
Theorem 3.3.
Under Assumption 3.2, the ML estimator is weak consistent,
$\hat{\Theta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Theta_{0}$
as $n\to\infty.$
Proof.
Here we assume that the covariate ${\bf X}$ is a continuous random vector with pdf $g$, but one can get the same result with a discrete random vector. Also, we assume $N$ is known, but the same result follows similarly if it lies in a known compact interval.
Note that the fractional binomial distribution is identifiable, since, by (3.1), we have $E(B_{N}(p,H,c))=Np,$ $P(B_{N}(p,H,c)=N)=p(p+c)^{N-1},$ and
$var(B_{N}(p,H,c))=Np(1-p)+\sum_{\begin{subarray}{c}i,j=1,2,\cdots,n\\ i\neq j\end{subarray}}pc|i-j|^{2H-2}.$
Therefore, $(p,H,c)$ uniquely determines the FB distribution, with which, combined with Assumption 3.2 ii), it follows that $\Theta$ uniquely determines the FB regression model, therefore, the identifiability of the FB regression model is established.
Under assumption 3.1, from Proposition 2.2 in [8] it follows that $P(B_{N}(p,H,c)=\ell)\neq 0$ for any $\ell=0,1,\cdots,N.$ Note that from (3.1-3.3) the log of the pmf of a fractional binomial distribution is continuous with respect to $(p,H,c^{\prime})$ which is, again, continuous with respect to the parameter $\Theta.$ Therefore, $\ln P(y_{i}|{\bf x}_{i},\Theta)$ is continuous in $\Theta$ for all $(y_{i},{\bf x}_{i}).$ Also, since $P(y_{i}|{\bf x}_{i},\Theta)$ is positive and continuous in $\Theta,{\bf x}_{i}$, by Assumption 3.2 i,iii, $|\ln P(y_{i}|{\bf x}_{i},\Theta)|<D(y_{i})<\infty$ for all $\Theta\in\mathcal{C},{\bf x}_{i}\in\mathcal{X}$ and $\int\sum_{y=0}^{N}P(y|{\bf x},\Theta_{0})D(y)dg(\bf{x})<\infty.$ Therefore, by Lemma 2.4 in [10]
$\sup_{\Theta\in\mathcal{C}}|\hat{\ell}(\Theta)-\ell(\Theta)|\stackrel{{\scriptstyle p}}{{\rightarrow}}0$
where
^$\hat{\ell}(\Theta|(y_{i},{\bf x}_{i}),i=1,2,\cdots,n)=\sum_{i=1}^{n}\ln P(y_{i}|{\bf x}_{i},\Theta)$
and
$\ell(\Theta)=\int\sum_{y=0}^{N}\ln P(y|{\bf x},\Theta)P(y|{\bf x},\Theta_{0})dg({\bf x}),$
and by Theorem 2.5 in [10]
$\hat{\Theta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Theta.$
Corollary 3.4.
The ML estimator of a fractional binomial distribution is weak consistent if the parameters lie in a known compact set, $(p,H,c^{\prime})\in\mathcal{D}$ where $\mathcal{D}$ is a compact set in $(0,1)^{3}.$
Note that the consistency of the ML estimator is held under the set of assumptions among which is that a compact set that includes the parameters is known. In the next section, through simulations, we check the performance and consistency of ML estimators that are obtained with a constraint of parameters in a known compact set. But, in practice, the assumptions often fail, especially as one may not know the possible range of parameters. Yet, the consistency of estimators is not of concern in data analysis as there is no guarantee that the data is generated from the assumed model. Rather the goal is to find a model that is most suitable for data. Therefore, when we fit the FB regression model to data in Section 5, we do not use a constraint of parameters to find MLEs.
4. Simulation
We simulate fractional binomial regression models generating $y_{i},i=1,2,\cdots,n=100,$ with covariates, ${\bf x}_{i}=(x_{i1},x_{i2}),$ i.i.d. uniform random variables from (-2,2), parameters $\Theta=(\psi_{1},\psi_{2},\eta_{1},\eta_{2},\nu_{1},\nu_{2})=(-1,1,2,1,0,-1),$ and $N=10.$ With the simulated data, $\{(y_{i},{\bf x}_{i}),i=1,2,\cdots,n\},$ we obtain MLE of the parameters with the constraint $\hat{\Theta}\in[-5,5]^{6}$. We repeat this procedure 20 times and compute the bias and standard error of the estimators. In the same way, the bias and standard error of the estimators are obtained with different values of $n,$ $\Theta$, and $N,$ and the results are shown in Table 1.
It is seen that overall, the estimates were fairly close to the true parameters as the bias and standard error are less than .5 in most cases. It is observed that, for a given set of parameters $\Theta$ and $N,$ if we increase the sample size $n,$ the standard error of each estimator decreases and the bias of each estimator also tends to decrease with few exceptions. This shows that the larger the sample size, the more precise the estimator, which was expected from the consistency of MLEs in Theorem 3.3.
N=10 N=20
n=100 n=400 n=100 n=400
Bias (s.e) Bias (s.e) Bias (s.e) Bias (s.e)
$\psi_{1}=-1$ $\hat{\psi}_{1}$01(.234) .003(.068) .064(.129) .004(.059)
$\psi_{2}=1$ $\hat{\psi}_{2}$027(.162) -.003(.059) .003(.096) .021(.058)
$\eta_{1}=2$ $\hat{\eta}_{1}$26(1.057) -.029(.33) -.035(.415) .011(.238)
$\eta_{2}=1$ $\hat{\eta}_{2}$209(.702) -.03(.206) -.074(.283) .064(.126)
$\nu_{1}=0$ $\hat{\nu}_{1}$225(1.068) .152(.261) .117(.567) -.031(.299)
$\nu_{2}=-1$ $\hat{\nu}_{2}$.305(.985) -.023(.268) -.355(.65) -.058(.308)
[…]
$\psi_{2}=2$ $\hat{\psi}_{2}$062(.266) .022(.074) .139(.262) .053(.112)
$\eta_{1}=1$ $\hat{\eta}_{1}$.049(.627) .018(.243) .026(.495) -.074(.179)
$\eta_{2}=-2$ $\hat{\eta}_{2}$.715(1.345) -.127(.367) -.44(.96) .013(.354)
$\nu_{1}=2$ $\hat{\nu}_{1}$66(1.474) .309(.753) .549(1.56) .291(.683)
$\nu_{2}=1$ $\hat{\nu}_{2}$.276(1.353) .3(.642) -.136(1.316) .258(.511)
Table 1. Bias and standard error of MLEs
5. Data analysis
5.1. Data from horticulture
In [11], zero-inflated models were fitted to a dataset that is from an experiment published in 1993 titled ”Micropropagation of columnar apple trees.” In their research, they micropropagated 270 shoots of the columnar apple cultivar Trajan. These shoots were treated with differing concentrations of cytokinin 6-benzylaminopurine (BAP). They were also treated under different photo periods; some were treated with an 8-hour photo period, whilst the rest were treated with a 16-hour photo period (Pho).
In this dataset, the response variable is the number of roots for each apple shoot, and BAP and photoperiod are two covariates. We fit the fractional binomial regression model and other zero-inflated regression models and compared their fit to the dataset. We fit each model with categorical variable BAP and numerical variable Pho, again with both variables treated as categorical.
Table 2 shows the AICs of the fitted models. The FB model has the lowest AIC among the zero-inflated regression models in both sets of covariates. Table 3 shows the results of Vuong’s test which is to see if the FB is significantly closer to the true generating process than other models to the true process. Its significance was found when compared to ZIP at a 5% significance level when BAP was numerical.
Figure 2 shows the probability distribution of the number of roots from the fitted models and from the data when categorical covariates were used with photoperiod 8, BAP 2.2 (left) and photoperiod 16, BAP 8.8 (right).
ZIP ZINB ZINB-2 FB
Pho(categorical) Log-likelihood -630.64 -621.95 -615.00 -611.48
BAP(numerical) AIC 1273.28 1257.9 1248 1240.96
Pho, BAP Log-likelihood -627.90 -620.03 -611.36 -606.32
(categorical) AIC 1275.80 1262.06 1252.72 1242.64
Table 2. Model comparison
test-statistics p-value
Pho(categorical) FB VS ZINB-2 0.971 0.166
BAP(numerical) FB VS ZINB 1.226 0.110
FB VS ZIP 1.723 .042
Pho, BAP FB VS ZINB-2 1.482 .069
(categorical) FB VS ZINB .508 .306
FB VS ZIP 1.144 .126
Table 3. Vuong’s closeness test between FB and other models
Figure 2. Estimated probability distributions and data distribution of the number of roots with Pho 6, BAP 2.2 (left) and Pho 16, BAP 8.8 (right).
5.2. Data from public health
The Ethiopian Mini Demographic Health Survey (EMDHS) collected data from a nationally representative sample on public health that includes the background characteristics of the respondents, the determinants of fertility, childhood mortality, maternal health care, etc. Among the variables is the number of antenatal care service visits by women age 15-49 who had a live birth in the 5 years before the survey. The data was analyzed in [5] with multi-level Hurdle negative binomial regression model.
We chose the number of antenatal care visits in the Oromia region of Ethiopia from the 2019 EMDHS, and analyzed its association with factors, age and educational level, through zero-inflated regression models.
The histogram of the number of antenatal care visits in Oromia is in Figure 3. It shows that 29.53% of the women in the sample (145 of 491 women) did not visit antenatal care.
Figure 3. Data distribution of the number of antenatal care visits
We used categorical covariates, age and educational level, and fit the fractional binomial regression model and other zero-inflated regression models. The estimation results are in Tables 4 and 5. When comparing ZIP, ZINB, and ZINB-2 regression models in Table 5, the additional terms in dispersion parameter $\theta$ in ZINB and ZINB-2 are not significant, and the estimates of each coefficient are very similar across the three models, which explains the fact that the log-likelihood of the three models appear to be the same in Table 6. It was found that the FB model has the lowest AIC, and when compared to other models through Vuong’s test, it is significantly better as shown in Table 7.
The probability distribution of the number of antenatal care visits by women aged 15-24 with primary education from the fitted models and its data distribution are shown in Figure 4 on the left. On the right of Figure 4, it shows the distributions of visits by women aged 25-34 with primary education. It is observed that the FB fits the data better at the peak in the middle and at the dip after 0 compared to other models.
logit link variable Categories Estimates (p-value) (Intercept) -1.450 (.000) Age 15-24 Ref 25-34 .211 (.043) $p$ 35-49 .046 (.759) Educational level Not educated Ref Primary .522 (.000) Secondary and higher .917 (.000) (Intercept) 11.725(.904) Age 15-24 Ref 25-34 1.368 (.992) $H$ 35-49 -1.683 (.986) Educational level Not educated Ref Primary .309 (.956) Secondary and higher -26.726 (.000) (Intercept) -1.666 (.000) Age 15-24 Ref 25-34 .054 (.805) $c^{\circ}$ 35-49 .355 (.150) Educational level Not educated Ref Primary -.460(.024) Secondary and higher -17.717 (.000)
Table 4. Fractional binomial regression with health data
log/logit variable Categories Estimates (p-value) link ZIP ZINB ZINB-2 (Intercept) 1.136 (.000) 1.136 (.000) 1.136 (.000) Age 15-24 Ref Ref Ref $\mu$ 25-34 .116 (.086) .117(.086) .116 (.087) (log) 35-49 .129 (.156) .129(.156) .129 (.157) Educational level Not educated Ref Ref Ref Primary .123 (.064) .123(.064) .123 (.064) Secondary and higher .170 (.067) .170(.066) .171 (.067) (Intercept) -.450(.098) -.450 (.098) -.448 (.100) Age 15-24 Ref Ref Ref $\pi$ 25-34 -.205 (.481) -.205(.481) -.206 (.479) (logit) 35-49 .358 (.265) .359(.264) .358 (.267) Educational level Not educated Ref Ref Ref Primary -.954 (.000) -.954 (.000) -.955 (.000) Secondary and higher -7.261 (.814) -8.317 (.897) -6.576 (.715) (Intercept) 17.413(.075) 13.404(.930) Age 15-24 Ref $\theta$ 25-34 3.482(.969) (log) 35-49 2.462(.923) Educational level Not educated Ref Primary 2.783(.977) Secondary and higher 2.245(.969)
Table 5. Zero-inflated regression models with health data
ZIP ZINB ZINB-2 FB Log-likelihood -886.13 -886.13 -886.13 -861.67 AIC 1793.26 1794.26 1802.26 1753.34
Table 6. Model comparison
test-statistics p-value FB VS ZINB-2 3.756 .000 FB VS ZINB 1.855 .032 FB VS ZIP 1.378 .084
Table 7. Vuong’s closeness test between FB and other models
Figure 4. Estimated probability distributions and data distributions of the number of antenatal care visits by women with primary education and aged 15-24(left) and aged 25-34(right)
6. Discussion and conclusion
We developed a new generalized linear model for count data that has many zeros. Our model utilizes the fractional binomial distribution that can serve as an over-dispersed, zero-inflated alternative to the regular binomial distribution. We used two datasets from agriculture and public health to fit zero-inflated Poisson/negative binomial regression and fractional binomial regression models. Our model has the lowest AIC in both datasets, and in some cases, it has a significantly better fit when compared to other models via Vuong’s test.
One limitation of the FB regression model is that it is not suitable for a large data set. The FB regression model is computationally expensive since the pmf of the fractional binomial distribution does not have a closed-form expression unlike Poisson and negative binomial distributions. This hinders it from fitting a large dataset, and it calls for scalable FB regression work for future work.